OpenStudy (anonymous):

The vector u = (3900, 5300) gives the number of units of two models of laptops produced by a company. The vector v = (2000, 1100) gives the prices (in dollars) of the two models of laptops, respectively. Identify the vector operation used to increase revenue by 5%.

OpenStudy (turingtest):

It's a scalar pruduct and a dot-product, right fewscrew?

OpenStudy (turingtest):

*product

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Yes.

OpenStudy (turingtest):

so it looks to me like\[1.05\overrightarrow{u}*\overrightarrow{v}\]where that is a dot product above

OpenStudy (turingtest):

the 1.05 should be tacked on to the v I guess, but it doesn't really matter as long as you don't distribute.

OpenStudy (turingtest):

So identifying the operation to increase revenue, it is a scalar multiplication by 1.05. The other operation is just to find the revenue in the first place. Seems to be what they want...

OpenStudy (anonymous):

so would it be u*(1.05)\[\left| v \right|\]

OpenStudy (anonymous):

or just (1.05)u*v

OpenStudy (anonymous):

wait i meant 1.05(u*v)

OpenStudy (anonymous):

if * represents the dot product, then they're all the same thing

OpenStudy (turingtest):

well it would be just the \[1.05\overrightarrow{v}\]to increase the price, and\[1.05\overrightarrow{v}*\overrightarrow{u}\]which is the same as you have. Just don't distribute the scalar to BOTHand it should be the same. ^Yeah you got it...

OpenStudy (anonymous):

@turingtest: that can't be distributed anyway\[c(\vec{u}\cdot\vec{v})\neq c\vec{u}\cdot c\vec{v}\]

OpenStudy (anonymous):

well the options are put in different ways : u*(1.05)\[\left| v \right|\] 1.05(u*v) 1.05\[\left| u*v \right|\] 1.05\[\left| u \right|\]*v

OpenStudy (turingtest):

That's what I am saying, note that there are no parentheses is my point.

OpenStudy (turingtest):

I'm trying to decipher what you wrote there ahaikal

OpenStudy (anonymous):

didnt you say the 1.05 had to be attached to the v as in U*(1.05)\[\left| v \right|\]

OpenStudy (anonymous):

yeah i know it got all broken up idk why

OpenStudy (anonymous):

its because i put that absolute value things

OpenStudy (turingtest):

it doesn't matter in the end because you can't distribute the 1.05 as pre-algebra pointed out. I am wondering if your bars are absolute value or magnitude, because you can't have the magnitude of a dot-product since it is already a number, and similarly you can't get the absolute value of a vector. So why are the symbols on both?

OpenStudy (anonymous):

typically, with |x|, if x is a scalar, then that's just the absolute value, and if x is a vector, then that's the length/magnitude of the vector.

OpenStudy (turingtest):

so if you take it that way^ then your formula is the same as mine

OpenStudy (turingtest):

absolute value isn't mentioned in mine because all values are positive.

OpenStudy (anonymous):

yes, but it must matter whether to put the parenthesis or not, or where to put the 1.05, because only one of the choices is correct. i think it is u*(1.05)v. (i just left our the absolute value sign or magnitude sign whichever)

OpenStudy (anonymous):

ahaikal10, like turingtest said, the only answer that makes sense here is 1.05(u*v)=1.05*u*v=u*1.05*v=u*v*1.05=etc.

OpenStudy (anonymous):

the parenthesis doesn't matter...

OpenStudy (turingtest):

so do you have a multiple choice and the things pre-algebra had are given as different choices?

OpenStudy (anonymous):

oh okay. yes i have multiple choice. so it would be 1.05(u*v)

OpenStudy (turingtest):

looks like it :)

OpenStudy (anonymous):

thank you very much guys :)